what is the dimensions for the rectangle that has the perimeter of 30 and the area of 56

Found 2 solutions by ikleyn, Shin123:
Answer by ikleyn(52926)

.
The sum of dimensions is


L + W = 30/2 = 15   and


the product is


L*W = 56


=============>  The dimensions are  7 units and 8 units.




Answer by Shin123(626)

L+w=30/2=15
l *w=56
7 *8=56
7+8=15
=========>Dimensions are 7 and 8.